OpenStudy (anonymous):

integrate (1/4)[(e^(4x))/(1+e^(4x))]e^(4x)

OpenStudy (adamaero):

\[(1/4)\ \int\limits_{}^{} \frac{ e ^{4x} }{ 1 + e ^{4x} }e ^{4x}dx\]

geerky42 (geerky42):

\[\int\dfrac{1}{4}\left[\dfrac{e^{4x}}{1+e^{4x}}\right]e^{4x}dx\]? You beat me @adamaero

OpenStudy (adamaero):

haha

OpenStudy (adamaero):

Have you tried u-sub?

OpenStudy (anonymous):

no I haven't I'm totally confused right now....

OpenStudy (adamaero):

combine the e^4x

OpenStudy (anonymous):

I know how to integrate but I just cant' think of how to use the u substitution, which one would be my u?

OpenStudy (perl):

u = 1 + 4e^x

OpenStudy (anonymous):

ok so du would be 1/4du

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

can someone explain the steps to me? I'm doing integrals also, don't really understand how du=1/4dx

OpenStudy (adamaero):

are you going to take this @eescobedo1 ?

OpenStudy (anonymous):

i think he left

OpenStudy (adamaero):